### v2

# Foundation Cures Personalization: Improving Personalized Models’ Prompt Consistency via Hidden Foundation Knowledge

Abstract
Facial personalization faces challenges to maintain identity fidelity without disrupting the foundation model’s prompt consistency. The mainstream personalization models employ identity embedding to integrate identity information within the cross-attention mechanisms of UNet. However, our preliminary experimental findings reveal that identity embeddings compromise the effectiveness of other tokens in the prompt, thereby limiting high prompt consistency and controllability. Moreover, by deactivating identity embedding, personalization models still demonstrate the underlying foundation models’ ability to control facial attributes precisely. It suggests that such foundation models’ knowledge can be leveraged to cure the ill-aligned prompt consistency of personalization models. Building upon these insights, we propose FreeCure, a framework that improves the prompt consistency of personalization models with their latent foundation models’ knowledge. First, by setting a dual inference paradigm with/without identity embedding, we identify attributes (e.g., hair, accessories, etc.) for enhancements. Second, we introduce a novel foundation-aware self-attention module, coupled with an inversion-based process to bring well-aligned attribute information to the personalization process. Our approach is training-free, and can effectively enhance a wide array of facial attributes in a non-intrusive manner; and it can be seamlessly integrated into existing popular personalization models, without harming their well-trained modules. FreeCure has demonstrated significant improvements in prompt consistency across a diverse set of state-of-the-art facial personalization models while maintaining the integrity of original identity fidelity. The project page is available here.
Figure 1: Personalization models (a) demonstrate strong capability in preserving identity fidelity, albeit at the cost of reduced prompt consistency. A prevalent feature in most personalization models is that when their identity embedding inputs are deactivated, they regain the ability to exhibit highly accurate prompt consistency with respect to facial attributes (b), a characteristic closely aligned to their foundation models. Our proposed FreeCure effectively leverages the latent foundational knowledge inherent in personalized models, enhancing prompt consistency in scenarios involving complex facial attribute control while preserving the identity fidelity (c).
1 Introduction
Human face-centric personalization represents compelling downstream applications of art creation, advertising, and entertainment in the realm of text-to-image synthesis  [17, 40]. Given a limited number of images that depict particular identities, facial personalization models generate novel content that reflects these identities through user-defined text prompts  [11, 35]. However, this aspiration is hindered by a persistent challenge: the necessity to maintain high fidelity to the identity while ensuring the controllability of the generated content, also referred to as prompt consistency  [53, 20]. This challenge is pronounced in facial personalization since any imperfections or misalignment in the generated faces are particularly salient to humans. Therefore, human face-centric personalization mandates dedicated attention and research efforts.
Previous research of facial personalization aims to integrate identity-specific information into the cross-attention mechanisms, with either fine-tuning based strategy  [35, 11] or tuning-free paradigm with an identity encoder  [46, 47, 27, 14]. However, the dual objectives of maintaining prompt consistency and identity fidelity remain unresolved. Despite the aforementioned advancements, state-of-the-art personalization techniques still struggle to enhance identity fidelity without sacrificing prompt consistency.
Our preliminary experiments on a wide range of personalization methods indicate that, under identical experimental conditions (e.g., prompts and initial random noise), personalization models’ outputs exhibit a significant decline in prompt consistency compared to generated results without identity embedding. For instance, as shown in the first two examples in Figure 1, personalization models fail to generate “blonde curly hair” and “angry” faces accurately, whereas their counterparts without identity embedding can handle these prompts in a highly faithful manner. Plus, when handling more complex prompts that consist of multiple attributes, the personalization model generates even poorer results. These findings show a fact overlooked by most previous works: while it is true that personalization models show their degradation in prompt consistency, the ability of their original foundation models is unharmed but overridden. Based on this observation, we step further to focus on the most critical components in personalization models: identity embedding and attention mechanisms. We find that identity embeddings can significantly undermine prompt consistency by impeding normal representation of other attribute-related tokens within the prompt through cross-attention mechanisms. This, in turn, adversely affects their effective expression in the latent space of the U-Net. However, given the compelling zero-shot identity extraction capability of identity embedding, directly modifying personalization models’ well-trained cross-attention modules can destroy their ability to capture precise identity information. Therefore, our core pursuit is clear: Is it feasible to mitigate the erosion of prompt consistency in personalization models while keeping their trained cross-attention modules unaffected?
Motivated by this objective, we propose FreeCure, a framework enhancing the prompt consistency of personalization models through the guidance of their latent foundation knowledge. While keeping cross-attention modules intact, we propose a novel foundation-aware self-attention (FASA), enabling attributes with high prompt consistency to replace those that are ill-aligned during personalization generation. To protect the identity unharmed, this strategy also leverages semantic segmentation models to generate the scaling masks of these attributes, therefore making such replacement happen in a highly localized and harmonious manner. Furthermore, we use a simple but effective approach called asymmetric prompt guidance (APG) to restore abstract attributes such as expression. Through comprehensive experiments, FreeCure has been verified to effectively restore a wide variety of misaligned attributes produced by various state-of-the-art personalization models.
In summary, the contributions of our paper are threefold:
  • We identify the limitations in prompt consistency prevalent across existing face-centric personalization models. Building upon this, we explore the negative effects of identity embedding and elucidate the fundamental reasons for its adverse impacts.
  • We propose a training-free framework that leverages high prompt consistency information from foundation models to enhance multiple weakened attributes in the synthesized results of personalization models. Crucially, the enhancement of various attributes is achieved without mutual interference.
  • Our framework can be seamlessly integrated into widely used personalization models and diverse foundational models. Comprehensive experiments show that our approach improves prompt consistency while maintaining the well-trained ability for identity preservation.
2 Related Work
Identity-preserving Generation. Identity-preserving generation can be broadly categorized into fine-tuning-based and encoder-based approaches. Fine-tuning-based methods  [11, 43, 1, 54, 10, 50, 31] either optimize a vector within a textual embedding to encode identity, modify specific weights in the U-Net [35, 23, 5, 19, 13, 36] or use low-rank decomposition techniques [18]. However, these methods require training a separate model for each identity, which limits their scalability and poses a risk of overfitting, particularly when working with limited data. In contrast, encoder-based methods [46, 48, 29, 27, 32, 7, 38, 26] focus on training an image encoder or a lightweight mapping network to derive identity representations aligning with textual embeddings. This allows for zero-shot personalization with novel identity references. Additionally, these methods also integrate key and value components of cross-attention maps into the training process [47, 32, 29, 44, 42, 39]. By employing detection and segmentation models [8, 21], they refine identity alignment and generation diversity by eliminating extraneous background elements. [14] integrates a lightning-fast T2I model to enable rapid sampling during personalization. Overall, encoder-based methods offer more flexible interactions between identity and textual embeddings but they still struggle to balance prompt consistency and identity fidelity.
Attention in Stable Diffusion. The attention mechanism is a fundamental element in Stable Diffusion.  [15, 2, 12, 22, 41, 4, 30, 9] have made attempts to edit important attention layers within diffusion models, allowing for localized attribute editing and style transfer.  [28] has explored the distinct roles of self-attention and cross-attention layers within the U-Net.  [24, 6, 3, 55] have investigated the latent properties of diffusion models to identify an editable semantic space via latent space analysis. With the advancement of inversion techniques  [40] in diffusion models, these methods aim to alter specific attributes of realistic images. Overall, the attention mechanism provides valuable insights into diffusion models. Nevertheless, within the realm of personalization, the role of identity embedding within the attention mechanism remains under-explored.
3 Revisit ID Embedding in Personalization
We conduct a comprehensive analysis to justify the limitations of current personalization methods in maintaining prompt consistency and to uncover the underlying factors contributing to this issue. This investigation not only highlights the challenges inherent in existing approaches but also provides a foundational basis and critical insights that inform the development of our proposed methodology.
Figure 2: Limitation of identity embedding in personalization models. (a) compares the impact of identity embeddings on the personalization and foundation outputs (zoom-in for better evaluation of prompt consistency). (b)-(d) compare scores of the cross-attention maps corresponding to identity embeddings and other attribute-related token embeddings.
3.1 Dual Denoising to Study Prompt Consistency
To elucidate current personalization methods’ limitations, we have devised a simple comparative experiment as an initial exploration. We maintain a fixed noisy latent code $z_{T}$ and implement two parallel denoising procedures. The only difference is that one denoising procedure incorporates textual embeddings fused with identity embedding $\tilde{c}$, while the other uses the original textual embeddings $c$, without any identity embedding information. For clarity, we will refer to the denoising procedure with identity embedding (personalization denoising) as PD, and the denoising process that excludes identity embedding (foundation denoising) as FD hereafter. These two procedures can be represented as the following equations:
$\textbf{PD}:\epsilon_{p}=\epsilon_{\theta}(z_{t},t,\tilde{c}),$ (1)
$\textbf{FD}:\epsilon_{f}=\epsilon_{\theta}(z_{t},t,c),$ (2)
where $\epsilon_{p}$ and $\epsilon_{f}$ denote the predicted noise from PD and FD, respectively. We have conducted these experiments using two facial personalization methods  [47, 27] and results are presented in Column (a) of Figure 2. It is evident that the output from PD exhibits a reduced ability to match the specific facial attributes compared to the counterpart without identity embedding. For instance: 1) The hair features do not align with the prompt’s specified “black curly” as expected; 2) The expression “laughing” is also restricted. Conversely, the results from FD show greater prompt consistency. These findings indicate that current identity embedding methods may not effectively address “the balance between identity fidelity and prompt consistency”.
3.2 ID Embedding’s Effect on Attention Layers
OBSERVATION 1: ID embeddings disrupt cross-attention layers. To investigate the underlying causes of this phenomenon, we conduct a visualization experiment utilizing the technique of cross-attention map analysis  [15, 2], because identity embedding is primarily fused into cross-attention layers. Columns (b-d) of Figure 2 present the visualization results of cross-attention maps for both PD and FD processes, focusing on identity embeddings and the attribute-specific token embeddings for “black curly hair” and “laughing”. It reveals that the identity embedding significantly amplifies activation in facial regions while simultaneously reducing activation for other tokens and disrupting their typical interactions within the cross-attention mechanism. This finding highlights identity embedding as a pivotal factor contributing to the reduced prompt consistency observed in personalization models. Additionally, we observe that during the FD process, where identity embedding injection is absent, prompt consistency remains preserved. This observation underscores a critical insight: although most personalization models are well fine-tuned, their inherent capability to generate faces with high prompt consistency is preserved but unexpectedly overridden by external identity embeddings. Thus, it is reasonable to conclude that the FD procedure functions in a manner analogous to pretrained foundation models  [34, 33], as implied by its full name, “Foundation Denoising”, and foundation models’ knowledge is still maintained in personalization models.
OBSERVATION 2: Cross-attention layers are highly susceptible. We conduct another experiment on cross-attention layers to prove its susceptibility, which involves incorporating a portion of the cross-attention map from FD’s placeholder token into the cross-attention map of PD’s identity embedding. The formula for this approach is
$A^{p}_{:,:,m}\leftarrow Softmax(\alpha A^{f}_{:,:,n}+(1-\alpha)A^{p}_{:,:,m}),$ (3)
where $A^{p}_{:,:,m}$ represents the cross-attention map of PD’s identity embedding, and $A^{f}_{:,:,n}$ represents the cross-attention map of the FD’s placeholder embedding (“man”, “woman”, etc.), whose positions in their respective prompts are $m$ and $n$. The parameter $\alpha$ controls the weight of the FD’s cross-attention map that is injected. Spatial indices $i$ and $j$ of $A$ are omitted since the interpolation operation is performed across the entire attention map. We apply this “interpolation” process to the same examples discussed in Figure 2 and the results are shown in Figure. 3. It is observed that, as the weight of the map from the FD process increases, the model incrementally loses identity fidelity, even though it regains the facial attributes that were missing before. This experiment further demonstrates that the cross-attention mechanisms in personalization models are rather susceptible to modifications. To preserve the models’ capacity for identity preservation, it is better to leave these well-trained cross-attention modules unaltered.
Figure 3: Interpolation of identity embedding’s cross-attention map. By injecting a portion $\alpha$ of the cross-attention map from the FD to the PD process, attributes can be restored. However, the identity fidelity can be easily undermined.
Figure 4: The overview of FreeCure. For a trained personalization model $\epsilon_{\theta}$, we first introduce (a): dual inference paradigm to generate faces with/without identity ($I_{p}$ and $I_{f}$), where $I_{f}$ without identity embedding shows better prompt consistency. Next, we leverage an external segmentation model $\Psi(\cdot)$ to derive related masks of target attributes with clear spatial information (hair, sunglasses, etc.) and merge them into a mask $\mathcal{M}$. In (b): we modify the original self-attention modules in the U-Net with our proposed FASA (c), which concatenates key and value matrices of FD process and PD process, together with a scaling mask to achieve the attribute injection. Finally, we utilize a simple yet effective strategy (d): asymmetric prompt guidance (APG) to restore abstract attributes (e.g., expressions).
OBSERVATION 3: Self-attention layers are the key to “foundation knowledge”. Our preliminary findings demonstrate that identity embeddings in cross-attention layers are disrupting personalization models’ prompt consistency. In contrast, its strong ability of identity extraction makes it rather challenging to be modified. Given this “dilemma”, we aim to explore a new approach to enhance prompt consistency by leveraging the self-attention layers in the U-Net. Since most personalization models introduce minimal modifications to self-attention layers, it is reasonable to assume that the aforementioned hidden foundation knowledge is preserved within them. By enhancing the self-attention layers in personalization models while keeping cross-attention layers intact, we anticipate achieving better alignment of facial attributes in personalized generation, all while minimizing adverse effects on identity fidelity.
4 Methodology
Building upon our investigation in Section 3, we propose FreeCure, a training-free framework designed to improve the prompt consistency of facial personalization models, particularly in contexts involving multiple attributes. We classify facial attributes into two distinct categories: those with well-defined spatial localization (e.g., hair, accessories) and those that are more abstract in nature (e.g., expression), implementing different strategies to enhance each category. Initially, we develop a foundation-aware self-attention (FASA) mechanism to integrate localized attributes from the FD process into the PD process (Section 4.1). Subsequently, we apply an asymmetric prompt guidance (APG) strategy to reconstruct more abstract attributes (Section 4.2). The overview is illustrated in Figure 4.
4.1 Foundation-Aware Self-Attention
Given a reference image $x_{ref}$ that provides the target identity, a well-trained personalization model $\epsilon_{\theta}$, a user-defined prompt $c$ that contains a sequence of facial attributes $\mathcal{A}=\{A_{1},A_{2},\cdot\cdot\cdot\,A_{n}\}$, our goal is to employ knowledge in FD to enhance prompt consistency of PD’s output.
Obtain Masks of Spatial Localized Attributes. As shown in Figure 4 (a), we adapt the dual inference branches which are identical to the FD/PD process introduced in Section 3 for personalization models to generate the personalized face $I_{p}$ with unsatisfactory prompt consistency as well as the foundation face $I_{f}$ with high prompt consistency. Next, we utilize an external face parsing model, denoted as $\Psi(\cdot)$, to extract the binary mask $M_{i}$ corresponding to $A_{i}$ from the foundation outputs $I_{f}$. Generally, when restoring multiple attributes, we simply need to compute the different masks respectively and merge them =$\mathcal{M}=\bigcup\{M_{i}\}$. $\mathcal{M}$ contains the spatial information of attributes that align with the target prompt, which will play an important role in the attribute restoration process that will be mentioned in the next part.
FASA Mechanism. FASA is the core module that links the information between FD and PD processes and enables personalization models to restore the attributes via their foundation knowledge. As shown in Figure 4(c), we at first identify the self-attention modules in the model. Specifically, in each timestep $t$ and attention layer $l$, we denote PD and FD’s key, query, and value matrices as $\mathcal{KQV}^{tl}_{p}=\{K^{tl}_{p},Q^{tl}_{p},V^{tl}_{p}\}$ and $\mathcal{KQV}^{tl}_{f}=\{K^{tl}_{f},Q^{tl}_{f},V^{tl}_{f}\}$. Second, we concatenate the key and value matrices of FD process to those of PD process: $\hat{K^{tl}}=[K^{tl}_{p},K^{tl}_{f}]$, $\hat{V^{tl}}=[V^{tl}_{p},V^{tl}_{f}]$. Thus, by omitting the labels of $t$ and $l$ for simplicity, the operation of FASA is
$\mbox{FASA}(\mathcal{KQV}_{p},\mathcal{KQV}_{f})=\mbox{Softmax}(\frac{Q_{p}\hat{K}^{T}}{\sqrt{d}})\hat{V}.$ (4)
Figure 5: Fine-grained attribute enhancement via masks. Extracting masks from the FD results makes the FASA module only focus on enhancement for target attributes, minimizing its negative effect on identity fidelity.
Fine-grained Restoration with Scaling Masks. The approach described in Equation 4 allows the PD process to obtain information from the FD process. However, we observe that the method also retains a substantial degree of unrelated features, leading to a notable loss of identity fidelity, as shown in Figure 5. To constrain attribute restoration in a specific region without disrupting identity information, we apply pre-calculated masks $\mathcal{M}$ of different attributes to the similarity map. Additionally, to further enhance FASA’s performance, we introduce an additional scaling factor, denoted as $\omega$, to control the magnitude of injecting attribute features from FD to PD. Therefore, the enhanced FASA mechanism can be written as
$\mbox{FASA}(\mathcal{KQV}_{p},\mathcal{KQV}_{f})=\mbox{Softmax}(\frac{[\mathbf{1},\omega\mathcal{M}]\odot Q_{p}\hat{K}^{T}}{\sqrt{d}})\hat{V}.$ (5)
Where $\mathbf{1}$ denotes a matrix with all elements equal to 1, designed to preserve attention on the original features within the PD process. $\odot$ represents the Hadamard product. With the enhanced FASA, attributes of the FD process can be successfully extracted and precisely injected into the PD process without disrupting the original personalization model’s ability of maintaining identity fidelity.
4.2 Asymmetric Prompt Guidance
Following attributes with clear location restored via FASA, to enhance more abstract attributes such as expressions, we introduce a simple yet effective method called asymmetric prompt guidance (APG). This strategy is based on the diffusion model’s inversion  [40]. During the inversion phase, the model accepts only a template prompt (e.g. “a man”) that does not include any target attributes. In the denoising process, we add the target attributes’ tokens back to this template prompt (e.g. “a man” $\rightarrow$ “a man laughing”). By leveraging the pretrained controllability of the foundation model, this approach enhances such attributes, resulting in the final refined image $I_{out}$. Throughout the denoising process, we use only pure textual prompts without identity embeddings, thereby avoiding their potential influence on the tokens related to the target attributes, as discussed in Section 3. Furthermore, to better preserve the identity, we start the denoising process directly from an intermediate latent code $\hat{z_{\gamma T}},\gamma\in[0,1]$, where the high-level identity information has already been established.
5 Experiments
Figure 6: Qualitative comparison with facial personalization baselines. Different attributes in prompts are highlighted in various colors. Comparison of corresponding FD outputs is provided in the Section 9 of the supplementary material.
Evaluation Datasets. We collect an extensive dataset encompassing both identity and prompt. The identity dataset comprises $50$ identities, with $30$ derived from the CelebA-HQ [25] and the other $20$ non-celebrity identities curated by our team. Each identity is represented by a single image, and the identities’ gender is balanced. Additionally, the identity dataset encompasses a spectrum of facial characteristics which better evaluates the stability of attribute enhancements. The prompt set consists of $20$ prompts: eight incorporate a single attribute, eight incorporate two attributes, and four incorporate three attributes. For each $<$identity, prompt$>$ pair, we produce 20 images for the evaluation process. Detailed information can be found in Section 8 of the supplementary material.
Baselines. We test FreeCure with five representative methods for facial personalization: PhotoMaker [27], PuLID [14], Face-diffuser [45], Face2Diffusion  [39] and FastComposer [47]. PuLID and PhotoMaker use SD-XL as their main foundation models, whereas Face-diffuser, Face2Diffusion, and FastComposer utilize SD v1-5. More details on how we set up each method’s dual branch of PD and FD can be found in the supplementary material.
Evaluation Metrics. We adopt CLIP-T [16] to calculate prompt consistency (PC). To calculate identity fidelity (IF), we use MTCNN [51] and FaceNet  [37] to extract the embedding of the generated/reference faces and compute the cosine similarity. Following PhotoMaker, we adopt the face diversity (Face Div.) metric which calculates LPIPS [52] between facial areas. Lastly, following Face2Diffusion, we adopt PC $\times$ IF score, since it reflects the overall balance of prompt consistency and identity fidelity. We compute the harmonic mean (hMean) of PC and IF.
FreeCure Settings. We set $\omega$ in FASA to $2.0$, to ensure that attribute information from FD can be sufficiently integrated into PD. The $\gamma$ in APG is set to $0.5$ to keep the balance of identity fidelity and prompt consistency restoration. For attribute segmentation, we leverage BiSeNet [49] and Segment-Anything [21] for different facial attributes.
5.1 Main Results
Overall Performance Comparison. Table 1 and Figure 6 present a comparative analysis of our proposed method with the baselines, examining both quantitative and qualitative aspects. Generally, the baseline methods often fail to accurately reflect key facial attributes mentioned in the prompts. For instance, attributes with large areas (hair, sunglasses, etc.) often fail to align with the target prompts. FastComposer, Face-Diffuser, and PuLID consistently generate faces with identical expression. This observation reinforces our earlier discussion in Section 3. Conversely, our approach shows a remarkable ability to bolster absent or faint attributes, significantly enhancing prompt consistency for these baselines. Secondly, we notice that incorporating FreeCure leads to a slight decline in identity fidelity. This can be attributed to the increased facial diversity in the enhanced results. In essence, since baseline personalization models tend to produce faces closely resembling their references, they inherently score higher in identity fidelity. However, this does not align with the objective of creating new content that adheres to diverse prompts. Ultimately, in terms of the PC $\times$ IF metric, our method also shows considerable improvement over all baselines. Qualitatively, the enhancements are particularly evident for localized attributes like hair and accessories and abstract attributes including expressions. In conclusion, both quantitatively and qualitatively, FreeCure demonstrates a positive balance by enhancing prompt consistency while keeping the reduction in identity fidelity to a minimum. Additional comparisons and results on more non-celebrity references are available in Section 9 of the supplementary material.
Method PC(%) $\uparrow$ IF(%) $\uparrow$ Face Div. (%) $\uparrow$ PC $\times$ IF (hMean) $\uparrow$
FastComposer (PD) 18.14 43.19 38.92 25.55
FastComposer + FreeCure 21.02 (+15.91%) 41.02 (-5.02%) 41.01(+5.37%) 27.80 (+8.82%)
Face-Diffuser (PD) 20.67 58.34 40.82 30.52
Face-Diffuser + FreeCure 22.48 (+8.76%) 57.51 (-1.42%) 41.95(+2.77%) 32.32 (+5.90%)
Face2diffusion (PD) 21.92 39.98 43.51 28.31
Face2Diffusion + FreeCure 23.26 (+6.12%) 39.23 (-1.88%) 44.29(+1.79%) 29.20 (+3.15%)
PhotoMaker (PD) 23.04 51.84 47.29 31.90
PhotoMaker + FreeCure 24.91 (+8.11%) 50.15 (-3.26%) 48.52 (+2.60%) 33.28 (+4.34%)
PuLID (PD) 25.16 58.23 42.12 35.14
PuLID + FreeCure 26.05 (+3.55%) 56.95 (-2.20%) 43.72(+3.80%) 35.74 (+1.74%)
Table 1: Main quantitative evaluation results. With FreeCure, the mainstream personalization models’ prompt consistency is highly enhanced on critical quantitative metrics.
Prompt Consistency with Multiple Attributes. Prompts involving multiple facial attributes pose a greater challenge to personalization’s prompt consistency but better reflect practical user needs. Table 2 illustrates FreeCure’s improvements on baselines with prompts including multiple facial attributes. As the number of attributes increases, the PC values of baselines tend to decrease. In contrast, FreeCure’s improvement in PC becomes more significant as the complexity of the prompt increases.
Method PC (1 Attr.) $\uparrow$ PC (2 Attr.) $\uparrow$ PC (3 Attr.) $\uparrow$
FastComposer (PD) 18.83 18.05 16.92
FastComposer + FreeCure 21.42 (+13.75%) 21.29 (+17.95%) 19.69 (+16.37%)
Face-Diffuser (PD) 21.68 20.98 18.02
Face-Diffuser + FreeCure 22.89 (+5.58%) 22.72 (+8.29%) 21.17 (+17.48%)
Face2Diffusion (PD) 22.54 21.98 20.54
Face2Diffusion + FreeCure 23.81 (+5.63%) 23.02 (+4.73%) 22.63 (+10.18%)
PhotoMaker (PD) 23.51 23.01 22.57
PhotoMaker + FreeCure 25.04 (+6.51%) 24.91 (+8.26%) 24.64 (+11.19%)
PuLID (PD) 25.56 25.12 24.43
PuLID + FreeCure 26.30 (+3.34%) 25.96 (+5.32%) 25.73 (+3.55%)
Table 2: Quantitative comparison of prompt consistency with different number of attributes.
Figure 7: Demonstration of FreeCure’s non-interference manner: top row: personalization models’ outputs; middle row: intermediate enhanced results with FASA; bottom row: final results after APG enhancement.
5.2 Robustness Justification
Non-interference Enhancement. We validate that the subsequent enhancement processes of FreeCure do not impact attributes that have already been enhanced. Figure 7 visualizes the intermediate results at FreeCure’s different stages. When APG is used in the latter stages, attributes previously enhanced through FASA remain unaffected. This illustrates that FreeCure’s improvement across multiple attributes is highly robust and consistent.
Robust Performance with Different Initial Noises. We observe that, with identical initial noise, all baselines’ FD and PD processes generate faces with similar attribute locations. It is important to validate that FASA’s robust performance when faces generated by FD and PD have different locations and angles. We relax this condition by using different initial noises and regenerating the faces. Figure 8 shows that results under the two settings are comparable, which confirms that FASA can effectively enhance the generated results of PD, even when its FD counterpart produces faces with variable spatial structures.
Figure 8: Performance of FASA w/ and w/o identical initial noises. FASA can precisely enhance attributes even if PD and FD produce faces with different locations, sizes, and angles.
Figure 9: Visualization of the FASA maps for attribute related area (red points) and non-attribute related area (green points).
Visualization of FASA maps. Figure 9 visualizes the FASA map $A\in\mathbb{R}^{H\times(2\times W)}$, whose size is doubled according to Equation 5. For a given query point $q_{i}$ in $Q_{p}$, its corresponding scores $A_{i,:}$ are extracted from both $K_{p}$ and $K_{f}$. When $q_{i}$ falls into areas associated with target attributes (e.g., hair, sunglasses), the FASA map exhibits greater attention to information from $K_{f}$, corresponding to FD. Conversely, in regions unrelated to attributes, the FASA map retains its attention on PD. This visualization substantiates the role of FASA in transferring fine-grained attribute information, equivalent to detailed foundational knowledge, from FD to PD. Additionally, FASA preserves its original attention pattern in regions unrelated to attributes, thereby ensuring the performance of personalization models in maintaining identity fidelity. More detailed analyses are available in Section 10 of the supplementary material.
5.3 Ablation Study
Overall Analysis of Each Component. Table 3 presents the individual performance of FASA and APG in enhancing attributes across five baselines. Both components demonstrate positive effects compared to the baseline metrics in Table 1, with FASA showing more noticeable improvements. We attribute this to FASA’s ability to effectively handle attributes covering larger areas, such as hair and sunglasses, resulting in more observable enhancements.
Method FASA APG PC(%) $\uparrow$ IF(%) $\uparrow$ PC $\times$ IF (hMean) $\uparrow$
FastComposer 20.91 41.94 27.91
19.03 42.84 26.35
Face-Diffuser 21.86 57.92 31.74
20.74 58.11 30.60
Face2Diffusion 22.73 39.30 28.80
21.98 39.47 28.24
PhotoMaker 24.76 50.62 33.25
23.61 50.86 32.25
PuLID 25.86 57.85 35.74
25.54 57.24 35.33
Table 3: Quantitative ablation analysis for FASA and APG’s independent effect on overall performance.
Figure 10: Ablation of scaling mask strategy.
Scaling Mask Strategy. Figure 10 illustrates the evaluation of the scaling mask strategy implemented in FASA. Low scaling values critically hinder FASA’s capacity to effectively transfer attribute information to the personalization denoising. On the contrary, high scaling factors may negatively impact the overall quality of the generated faces. Generally, the optimal value of $\omega$ should be around $2.0$. More ablation studies of FASA are available in Section 10.3 of the supplementary material.
Figure 11: Ablation of APG with different intermediate timesteps.
Effect of Inversion’s Intermediate Timesteps. Figure 11 demonstrates the effects of different starting timesteps in APG, represented by the parameter $\gamma$ in Section 4.2. To achieve a noticeable improvement in attributes while maintaining identity fidelity, an optimal $\gamma$ should be set as $0.5$.
6 Conclusion
Facing the challenge that personalization models employing identity embeddings frequently struggle to preserve prompt consistency, we introduce FreeCure, a training-free framework that leverages the high prompt consistency inherent in foundation models to refine the output of personalization models, leading to a remarkable improvement in their prompt consistency. Additionally, FreeCure is able to enhance multiple attributes concurrently in a manner that is both complementary and non-disruptive. Our experiments validate the effectiveness of FreeCure on popular baselines particularly in scenarios with complex prompts that encompass multiple attributes.
Supplementary Material
7 Overview
The Supplementary Material consists of three sections, providing additional details and results to complement the main submission.
  • Section 8 provides implementation details, including details of setting up the dual denoising paradigm for the FD and PD processes, segmentation models as well as evaluation metrics.
  • Section 9 provides additional results, including complementary FD outputs corresponding to the results included in Figure 6 of the main paper, additional experimental results on non-celebrity reference images and broader attribute enhancement achieved through the integration of advanced models, such as Segment-Anything.
  • Section 10 provides a comprehensive analysis of FreeCure, including analysis for initial noise conditions, detailed visualization of FASA modules, additional ablation studies to determine the optimal placement of FASA within the U-Net, and additional investigations into identity embeddings.
8 More Implementation Details
8.1 Triggering Hidden Foundation Knowledge in Personalization Baselines
The core architecture of FreeCure is based on a dual denoising paradigm of the FD and PD process, which triggers hidden foundation knowledge in personalization models. This paradigm deactivates the identity embeddings which are integrated into the cross-attention layers. In Table 4 we illustrate how we deactivate identity embedding for different baselines mentioned in the main paper since they have various identity embedding fusion strategies and prompt templates.
Method PD inputs FD inputs
FastComposer a <P> <|image|> with black curly hair a <P> image with black curly hair
Face-Diffuser a <P> <|image|> with black curly hair a <P> image with black curly hair
Face2Diffusion a f l with black curly hair a <P> image with black curly hair
PhotoMaker a <P> img with black curly hair a <P> image with black curly hair
Table 4: Strategies for implementing dual denoising paradigms on different baselines. “black curly hair” is used here as an example and can be replaced with other prompts containing different facial attributes.
FastComposer and Face-Diffuser share the same prompting template, wherein their identity embeddings are integrated into the “<P>” word, which is initialized with class-specific words (e.g., man, woman, boy, girl). To set FD process, we keep this “<P>” as its original plain text’s embedding without integrating any identity information. In the case of PuLID, which utilizes independent residual cross-attention adapters for identity fusion, we initialize its identity embedding as a zero tensor of the same dimensionality.
8.2 Attribute Segmentation Models
As mentioned in Section 5, we use two mainstream segmentation models: BiSeNet and Segment-Anything models for attribute-aware mask generation. BiSeNet  [^1] https://github.com/CoinCheung/BiSeNet is a popular framework for face parsing that is capable of generating semantic masks corresponding to facial attributes. Table 5 shows BiSeNet’s prediction label and corresponding facial attributes. BiSeNet can address the majority of facial attributes, thereby verifying the robustness of FreeCure. Additionally, FreeCure is designed to be seamlessly integrated with various segmentation models, not limited to BiSeNet. To validate this flexibility, we have showcased results that utilize Segment-Anything as the face parsing model in Section 9.3.
label attribute label attribute label attribute
0 background 6 glasses 12 lower lip
1 facial skin 7 right ear 13 neck
2 right eyebrow 8 left ear 14 necklace
3 left eyebrow 9 nose 15 cloth
4 left eye 10 earrings 16 hair
5 right eye 11 upper lip 17 hat
Table 5: Information for facial attribute and label relationship of BiSeNet. All critical attributes for the human face are included.
8.3 Details of Metrics
Prompt Consistency (PC, also known as CLIP-T). We leverage the official implementation of the vision transformer [^2] https://github.com/openai/CLIP provided by OpenAI. We utilize its image and text encoder to derive image and textual embeddings respectively. Ultimately, the cosine similarity between these two embeddings is computed as a quantitative measure of prompt consistency.
Identity Fidelity (IF). We use the official implementation of MTCNN + FaceNet pipeline [^3] https://github.com/timesler/facenet-pytorch to conduct the processes of face detection and feature extraction from facial regions. Additionally, we compute the cosine similarity between two face embeddings to evaluate their similarity.
Face Diversity (Face Div.). We utilize the official implementation of LPIPS [^4] https://github.com/richzhang/PerceptualSimilarity to quantify the perceptual distance between two facial images.
8.4 Facial Prompts for Evaluation
Table 6 introduces the prompts for evaluating enhancement performance for facial attributes. Generally, our prompts include multiple facial attributes, ensuring that previous baselines’ weaknesses in prompt consistency can be fully uncovered and highlighting the enhanced performance of FreeCure.
Attribute Prompt
1 a <S> with black curly hair
a <S> with blonde curly hair
a <S> with red long straight hair
a <S> with very angry looking
a <S> with frowning worriedly
a <S> laughing happily
a <S> wearing a mask
a <S> wearing sunglasses
2 a <S> with white curly hair, frowning worriedly
a <S> with black curly hair, laughing happily
a <S> with blonde curly hair and blue eyes
a <S> with blue eyes, laughing happily
a <S> wearing sunglasses, laughing happily
a <S> with black hair, wearing silver earrings
a <S> with blonde hair and blue eyes
a <S> with sunglasses and silver earrings
3 a <S> with red curly hair, wearing pearl earrings, unhappy looking
a <S> with blue eyes and blonde curly hair, smiling
a <S> with white curly hair, wearing sunglasses, laughing happily
a <S> with black curly hair, wearing silver earrings, frowning worriedly
Table 6: Prompts for evaluation categorized by the number of included attributes (range from 1 to 3). <S> will be replaced with placeholder tokens such as man, woman, boy, etc.
9 More Results
9.1 Supplementary Comparison of Results for Main Paper
To better demonstrate the efficacy of FreeCure, Figure 12 presents a comparative analysis of the results generated by the FD, PD, and enhanced outputs by FreeCure linked to Figure 6. The results indicate that FreeCure faithfully transfers target prompt-aligned attributes from the FD into the PD processes while preserving the original identity information.
9.2 Results on Non-Celebrity References
In Figure 13, we provide additional experimental results on non-celebrity reference images to further validate the consistent performance of FreeCure. This robust performance shows the potential of FreeCure to enhance various personalization models in practical applications, as the ability to handle non-celebrity personalization is a critical requirement in real-world scenarios.
9.3 Integration with Segment-Anything for Attribute Extraction
Figure 14 showcases FreeCure’s consistent performance when integrated with more advanced segmentation models such as Segment-Anything. For instance, the attribute mask cannot be handled by BiSeNet since its label is absent. However, by replacing BiSeNet with Segment-Anything and specifying the target prompt as “mask,” valid semantic masks can be accurately generated, thereby guaranteeing effective attribute enhancement through FASA. Given that Segment-Anything extracts attributes based on flexible prompts, its seamless integration significantly enhances the versatility of FreeCure, enabling it to address a broader range of scenarios.
10 More Analysis
10.1 More Comparison Results w/ and w/o Identical Initial Noise
We conduct extensive experiments to evaluate the impact of initial noise conditions. Figure 15 demonstrates that under identical initial noise conditions, FreeCure consistently enhances target attributes across different input reference images. This finding highlights the robustness and practical applicability of FreeCure. Figure 16 illustrates that FreeCure can successfully inject attribute information from the FD process to the PD process even if their final generated faces have various locations, angles, and sizes. This finding particularly demonstrates the robustness of the FASA module.
10.2 More Visualization of FASA
We provide a more detailed visualization of FASA in Figure 17, which substantiates the claim that FASA enhances facial attributes in a fine-grained manner. For instance, FASA effectively captures and faithfully transfers information of attributes localized in small areas (e.g. eyes, pearl earrings). Furthermore, in regions unrelated to the target attributes, FASA maintains a strong alignment with the original PD attention maps, demonstrating that it preserves the core functionality of the personalization models.
10.3 Ablations of Applying FASA at Different Positions in the U-Net
We conduct detailed ablation studies to determine the optimal placement of the FASA module within the U-Net, as illustrated in Figure 18. When FASA is applied exclusively to the downsampling and middle blocks, its effectiveness is limited. In contrast, applying FASA to the upsampling blocks yields the most significant performance improvements. Furthermore, adding FASA to the middle and downsampling blocks provides negligible additional benefits when the upsampling blocks are already utilized, as the upsampling blocks alone are sufficient to accurately transfer attribute information from the FD to the PD process. Based on these findings, we conclude that applying FASA to the U-Net’s upsampling blocks represents the optimal configuration.
10.4 More Studies on Identity Embedding
Figure 19 illustrates the impact of identity embedding interpolation when integrated into the cross-attention layers of different blocks, as mentioned in Section 3. When cross-attention maps of the FD process are injected into the downsampling blocks or middle blocks, the changes in output are minimal, even if all identity embedding’s cross-attention maps are replaced ($\alpha=1$). It is only when the interpolation is applied to the upsampling blocks that significant degradation of identity information is evident, while other facial attributes are effectively restored. In summary, the identity embedding exerts its most significant influence within the upsampling blocks of the U-Net. However, a small value of $\alpha$ can cause significant identity information loss, supporting the argument in Section 3, that the well-trained cross-attention layers for identity information extraction is susceptible.
Figure 12: Corresponding comparative FD & PD outputs for results in Figure 6 in the main paper. (a) refers to original personalization outputs, (b) refers to foundation outputs, and (c) refers to outputs enhanced with proposed FreeCure.
Figure 13: More results for non-celebrity references.
Figure 14: Results for FreeCure’s integration with Segment-Anything. FreeCure can seamlessly integrate with more advanced segmentation models such as Segment-Anything for extraction of a boarder range of attributes. Target prompt: “a person wearing a mask”
Figure 15: More studies for attribute enhancement start with identical initial noise. FreeCure exhibits robust performance with the single foundation output’s guidance on personalized outputs from various references (generated based on Face-Diffuser and PhotoMaker).
Figure 16: More validation on FreeCure’s robust performance with different initial noises for PD and FD processes. Even if PD and FD’s outputting faces have totally different locations, angles, and sizes, FreeCure can still exhibit stable enhancement performance.
Figure 17: More visualization of FASA map on attribute-relevant and attribute-irrelevant areas. Red points refer to the area of target attributes and green points refer to the area which is not related to target attributes.
Figure 18: Ablation study for applying FASA at different positions of UNet. We find out that applying FASA modules to upsampling blocks alone will be sufficient enough to exhibit promising enhancement for target attributes.
Figure 19: Identity embedding interpolation’s influence on cross-attention layers of different blocks in the U-Net. It is applied on: downsampling blocks (top), bottleneck blocks (middle), and upsampling blocks (bottom).